# Unveiling the Role of ChatGPT in Software Development: Insights from Developer-ChatGPT Interactions on GitHub

### v2

@@ 3.5. Data Extraction and Analysis
In this work, we examined how developers leverage ChatGPT throughout the software development process. For answering RQ1, we presented the distributions and characteristics of the shared ChatGPT links on GitHub, and we further analyzed the contents of shared ChatGPT conversations to answer the rest RQs (i.e., RQ2, RQ3, and RQ4). For the quantitative analysis addressing characteristics (RQ1) of shared ChatGPT links, we applied descriptive statistics. Concretely, we computed frequencies and percentages of the shared ChatGPT links, calculated the central tendency (mean, median, and mode) and dispersion (including the interquartile range), and visualized results distributions with histograms and bar plots. To examine and classify developers’ purposes (RQ2), activities (RQ3), and tasks (RQ4) when sharing ChatGPT links, we adopted the Constant Comparison method (Charmaz, 2014; Stol et al., 2016), a grounded theory technique rigorously validated for deriving emergent categories from qualitative data in software engineering research. Constant Comparison was chosen specifically because it is uniquely suited for exploratory qualitative research where categories (e.g., developers’ purposes) must emerge inductively from unstructured data, rather than being imposed a priori. According to the description provided in (Charmaz, 2014), the Constant Comparison process consists of three main steps: (1) initial coding, three authors (the first, third, and fourth authors) independently reviewed the content of shared ChatGPT conversations. For example, we initially labeled 22 purposes for RQ2 before the subsequent label merge (i.e., focused coding). (2) focused coding, the three authors selected the most frequent codes and categorized the data accordingly, with the results subsequently reviewed by the first author. For instance, we consolidated codes such as “Dataset Split”, “Data Merge”, and “Data Cleaning” into “Data Handling”, and merged “Program Comprehension” into “Code Explanation and Understanding”. (3) theoretical coding, the three authors aimed to establish relationships between the identified codes. For instance, we grouped the “Lifecycle Activities” and “Supporting Activities” into the root node “Development Related Activities” for RQ3. To minimize personal bias, disagreements in coding results were resolved through collaborative, multi-turn discussions among all the authors, ensuring an objective interpretation of the data. Note that, we extracted and stored relevant data entities from shared ChatGPT links in MS Excel files (see the dataset at (Li et al., 2025)) to facilitate data analysis. Our dataset also lends itself to further analysis of other topics, such as prompts, responses, and relevant interactions.
@@ 4.1. RQ1: Characteristics of ChatGPT Usages
Notably, Code emerges as the dominant share of interactions, comprising 43.4% (1,105) of all shared URLs. This reflects that developers most frequently use ChatGPT in contexts directly related to source code, such as generating or modifying code snippets. Commits follow as the second largest share at 32.3% (823), indicating that shared ChatGPT links are commonly referenced during the commit process. The remaining sources show comparatively lower usage: Issues account for 11.6% (295), Pull Requests for 10.4% (266), Discussions for only 2.3% (58), reflecting limited reliance on ChatGPT for open-ended or collaborative discussions. Overall, the distribution indicates that developer-ChatGPT conversations are most frequently shared in contexts directly related to code development and integration activities, while its presence in collaborative or discussion-oriented contexts is comparatively lower.
@@
Our DevChat dataset also contains the prompt data of developer-ChatGPT interactions. In this context, a turn refers to one complete prompt-response pair between a developer and ChatGPT (i.e., one developer’s question followed by one ChatGPT’s response). Figure 5 provides a box plot visualization of the prompt turns distribution across five GitHub sources, after filtering outliers. The corresponding statistics are summarized in Table 3, which includes the first (Q1), second (Q2, a.k.a Median), third (Q3) quartiles of the distributions, and the InterQuartile Range (IQR), along with the mode and its corresponding frequency. Together, Figure 5 enables a rapid visual assessment of the prompt turns distribution, and Table 3 provides a precise quantitative comparison of the prompt turn numbers.
@@
RQ1 Summary: (1) The shared ChatGPT links are predominantly from Code (43.4%) and Commits (32.3%), and the usage of the shared links peaked in August 2023, with consistently high engagement in Code and Commits. (2) ChatGPT links are most frequently shared in code-related tasks like code generation. The prompt turns distribution suggests that developers prefer short and task-focused interactions (e.g., concept interpretation, code modification). (3) Most developers include contextual descriptions when sharing ChatGPT links, particularly in Commits, Discussions, and Pull Requests, but the shared ChatGPT links in Code often omit contextual descriptions.
4.2. RQ2: Developers’ Purposes for Sharing ChatGPT Conversations
To answer RQ2, we employed the Constant Comparison method (Charmaz, 2014; Stol et al., 2016) to analyze and categorize the developers’ purposes for sharing ChatGPT links during software development. Specifically, we examined the textual context surrounding the shared links, including the link description, following messages, and explicit user intent statements; for those links without descriptions, we analyzed their developer-ChatGPT conversations. We finally got five primary categories of developers’ purposes (i.e., Task Delegation, Problem Resolution, Knowledge Acquisition, Solution Recommendation, and Concept Interpretation), and we analyzed their prevalence across the five sources. Figure 7 presents the distribution of the five purposes across five sources. Each circle is scaled to the total number of shared ChatGPT links for different sources, and the segments within each circle show the relative contribution of each purpose, making cross-category comparisons intuitive. For each category, we provide detailed descriptions of the results, followed by one example from our curated dataset. Detailed results are presented below.
Figure 7. Developers’ purposes for sharing ChatGPT links from the five sources
@@ 4.2. RQ2: Developers’ Purposes for Sharing ChatGPT Conversations
RQ2 Summary: We identified and categorized developers’ purposes for sharing ChatGPT links during software development into five categories. The results show that Task Delegation is the dominant purpose (75-88% Code/Commits/Pull Requests), primarily for automating repetitive tasks. Problem Resolution (8-11% in Issues/Commits) and Concept Interpretation (¡6%) indicate GenAI’s context limitations. Knowledge Acquisition (21% in Discussions) and Solution Recommendation (17% in Discussions) thrive in collaborative ideation.
4.3. RQ3: Development-Related Activities Involved in Shared ChatGPT Conversations
To address RQ3, which focuses on the development-related activities where developers share ChatGPT conversations, we analyzed the distribution of shared ChatGPT links across the five data sources. Specifically, we applied the Constant Comparison method (Charmaz, 2014; Stol et al., 2016) to establish a categorization of the development-related activities, considering all extracted data items including filenames, path names, and other metadata (see Section 3.5). For example, requirements and user stories are categorized as “Requirement Analysis”; explicit change requests and regression fixes discussions are categorized as “Software Maintenance and Evolution”; and path names containing “xxx/testing” or mentions of unit tests are classified as “Software Testing”. Besides, software documentation includes various documents which can be integrated in lifecycle activities (e.g., requirements specifications, architecture and design documents, test reports). Considering that shared ChatGPT links also serve as a form of documentation recording developer-ChatGPT interactions during software development, we chose not to treat software documentation as a separate activity. Otherwise, all the shared ChatGPT links would fall under this documentation activity. The results of RQ3 reveal the breadth of developer interactions potentially facilitated by ChatGPT, as illustrated in Figures 8,  9, and 10.
@@ 4.3. RQ3: Development-Related Activities Involved in Shared ChatGPT Conversations
Although there are various ways to categorize development-related activities, we considered and adapted the classification proposed in SWEBOK (Washizaki et al., 2024) and previous studies (Hou et al., 2024; Fan et al., 2023), which are widely recognized in the field of software engineering. Following this, we categorized the relevant activities involving shared developer-ChatGPT conversations, as shown in Figure 8. The development-related activities were then grouped into two primary categories: Lifecycle Activities and Supporting Activities.
  • Lifecycle Activities encompass seven conventional phases of the software development lifecycle (SDLC), that is, Requirements Analysis, Software Design, Software Implementation, Software Testing, Software Deployment, Software Maintenance and Evolution, and Software Management.
@@
RQ3 Summary: Analysis of shared ChatGPT links reveals that: (1) The number of shared ChatGPT links in the two lifecycle activities (i.e., Software Implementation and Software Maintenance & Evolution) far exceeds the number of other seven activities; (2) Software Implementation, Software Maintenance & Evolution are the dominant activities in which developers frequently share ChatGPT links.
4.4. RQ4: Tasks of ChatGPT Conversations Involved in Software Development
To answer RQ4, we investigated the specific tasks within developer-ChatGPT interactions during development. Here, we adopted a broader view of tasks by conducting an in-depth analysis of developers’ intents, which encompass both action-oriented tasks (e.g., code generation) and knowledge-oriented goals (e.g., understanding specialized tools and techniques). While the latter may not constitute a task in the strict operational sense, we considered them equally relevant for characterizing developers’ practical use of ChatGPT, since seeking such knowledge is often an essential step toward completing a development task. Figure 11 presents a Sankey diagram that visually decomposes nine development-related activities (see Section 4.3) into fine-grained SE tasks. This visualization highlights the relationships between the five data sources, nine development-related activities, and 39 detailed SE tasks derived from ChatGPT interactions in our collected GitHub data, thereby reflecting how developers utilize ChatGPT in their daily workflows.
@@ 4.4. RQ4: Tasks of ChatGPT Conversations Involved in Software Development
RQ4 Summary: Our analysis revealed 39 specific SE tasks across nine development-related activities, with Code Generation & Completion and Code Modification & Optimization emerging as the most prevalent SE tasks during software development. While code generation and maintenance tasks dominate, significant engagements were also observed in other crucial activities, highlighting ChatGPT’s broad utility as an assistant throughout the SE lifecycle.
@@ 5.1.1. Characteristics of ChatGPT Usage on GitHub
5.1.2. Developers’ Purposes for Sharing ChatGPT Conversations
@@ 5.2. Implications
For researchers, the findings point to promising opportunities to extend ChatGPT’s utility in certain areas, such as requirements engineering, by tailoring LLMs to domain-specific SE workflows. For example, the concentration of SE tasks in Software Implementation and Software Maintenance & Evolution implies that ChatGPT is primarily perceived as a coding assistant. Moreover, the findings highlight research opportunities to integrate ChatGPT in certain tasks (e.g., System Architecture Design), thereby improving its capability to boost productivity in underexploited areas. Besides, the study underscores ChatGPT’s transformative potential - not only in reshaping how developers engage in SE tasks, but also in redefining what constitutes a development-related activity in the AI era. This evolution calls for foundational knowledge in software engineering, such as SWEBOK (Washizaki et al., 2024), to adapt alongside emerging GenAI-driven software development practices.
@@
While this study provides valuable insights regarding the daily use of ChatGPT during software development, future research could delve deeper into the impact of ChatGPT on developer productivity. This could involve conducting controlled experiments to compare outcomes with and without ChatGPT’s assistance.
For practitioners, understanding the impact of ChatGPT is crucial for strategic technology adoption during development. Although GenAI tools like ChatGPT exhibit great potential in enhancing developer efficiency (Vaillant et al., 2024), their limitations, such as the propensity to generate inaccurate or misleading information (often referred to as hallucination (Achiam et al., 2023)) or even unsafe code content (Chen et al., 2025), cannot be overlooked. Therefore, for researchers, this necessitates the collection of robust empirical evidence to assess the reliability and practical utility of ChatGPT in SE tasks. Future investigations should focus on determining the contexts in which these tools deliver substantial benefits and the contexts where they may introduce risks or inefficiencies (Yu et al., 2024). Addressing these concerns will be critical to ensuring the responsible and effective integration of LLMs into modern software development workflows.
@@
Our results reveal that developers utilize ChatGPT in various development-related activities (particularly Software Implementation and Software Maintenance & Evolution). However, errors in ChatGPT’s outputs could potentially introduce inefficiencies or vulnerabilities in software systems (Liu et al., 2024b, a). Moreover, excessive cognitive reliance on GenAI tools might hinder the development of essential manual problem-solving skills, especially for junior developers. Given the potential of errors and biases in LLM-generated content (Chen et al., 2025; Li et al., 2024), it is imperative to establish robust safeguards, such as human oversight, rigorous validation processes, and strict adherence to ethical and legal standards. These measures are critical to maintaining the quality, security, and trustworthiness of AI-assisted development.

### v4

@@ 3.5. Data Extraction and Analysis
In this work, we examined how developers leverage ChatGPT throughout the software development process. For answering RQ1, we presented the distributional and usage characteristics of the shared ChatGPT links on GitHub, and we further analyzed the contents of shared ChatGPT conversations to answer the rest RQs (i.e., RQ2, RQ3, and RQ4). For the quantitative analysis addressing characteristics (RQ1) of shared ChatGPT links, we applied descriptive statistics. Concretely, we computed frequencies and percentages of the shared ChatGPT links, calculated the central tendency (mean, median, and mode) and dispersion (including the interquartile range), and visualized results distributions with histograms and bar plots. To examine and classify developers’ purposes (RQ2), activities (RQ3), and tasks (RQ4) when sharing ChatGPT links, we adopted the Constant Comparison method (Charmaz, 2014; Stol et al., 2016), a grounded theory technique rigorously validated for deriving emergent categories from qualitative data in software engineering research. Constant Comparison was chosen specifically because it is uniquely suited for exploratory qualitative research where categories (e.g., developers’ purposes) must emerge inductively from unstructured data, rather than being imposed a priori. According to the description provided in (Charmaz, 2014), the Constant Comparison process consists of three main steps: (1) initial coding, three authors (the first, third, and fourth authors) independently reviewed the content of shared ChatGPT conversations. For example, we initially labeled 22 purposes for RQ2 before the subsequent label merge (i.e., focused coding). (2) focused coding, the three authors selected the most frequent codes and categorized the data accordingly, with the results subsequently reviewed by the first author. For instance, we consolidated codes such as “Dataset Split”, “Data Merge”, and “Data Cleaning” into “Data Handling”, and merged “Program Comprehension” into “Code Explanation and Understanding”. (3) theoretical coding, the three authors aimed to establish relationships between the identified codes. For instance, we grouped the “Lifecycle Activities” and “Supporting Activities” into the root node “Development Related Activities” for RQ3. To minimize personal bias, disagreements in coding results were resolved through collaborative, multi-turn discussions among all the authors, ensuring an objective interpretation of the data. Note that, we extracted and stored relevant data entities from shared ChatGPT links in MS Excel files (see the dataset at (Li et al., 2025)) to facilitate data analysis. Our dataset also lends itself to further analysis of other topics, such as prompts, responses, and relevant interactions.
@@ 4.1. RQ1: Characteristics of ChatGPT Usages
Notably, Code emerges as the dominant share of interactions, comprising 43.4% (1,105) of all shared URLs. We can observe that developers most frequently use ChatGPT in code contexts, though the specific purposes require further analysis of the conversation content. Commits follow as the second largest share at 32.3% (823), reflecting that shared ChatGPT links are commonly referenced during the commit process. The remaining sources show comparatively lower usage: Issues account for 11.6% (295), Pull Requests for 10.4% (266), Discussions for only 2.3% (58), indicating limited reliance on ChatGPT for open-ended or collaborative discussions. Overall, the distribution shows that developer-ChatGPT conversations appear less often in Discussions than in Code and Commits; this may reflect different practices regarding how ChatGPT links are shared across GitHub sources.
@@
Our DevChat dataset also contains the prompt data of developer-ChatGPT interactions. In this context, a turn refers to one complete prompt-response pair between a developer and ChatGPT (i.e., one developer’s question followed by one ChatGPT’s response). Figure 5 provides a box plot visualization of the prompt turns distribution across five GitHub sources. Note that outliers (i.e., a few prompt turns over 100, see Table 3) are not shown to ensure an appropriate scale for the coordinate axes. The corresponding statistics are summarized in Table 3, which includes the first (Q1), second (Q2, a.k.a Median), third (Q3) quartiles of the distributions, and the InterQuartile Range (IQR), along with the mode and its corresponding frequency. Together, Figure 5 enables a rapid visual assessment of the prompt turns distribution, and Table 3 provides a precise quantitative comparison of the prompt turn numbers.
@@
RQ1 Summary: (1) The shared ChatGPT links are predominantly from Code (43.4%) and Commits (32.3%), and the usage of the shared links peaked in August 2023, with consistently high engagement in Code and Commits. (2) The prompt turns distribution of shared ChatGPT links suggests that developers prefer short and task-focused interactions. (3) Most developers include contextual descriptions when sharing ChatGPT links, particularly in Commits, Discussions, and Pull Requests, but the shared ChatGPT links in Code often omit contextual descriptions.
4.2. RQ2: Developers’ Purposes of Using ChatGPT Conversations
To answer RQ2, we employed the Constant Comparison method (Charmaz, 2014; Stol et al., 2016) to analyze and categorize the developers’ purposes of using ChatGPT links during software development. Specifically, we examined the textual context surrounding the shared links, including the link description, following messages, and explicit user intent statements; for those links without descriptions, we analyzed their developer-ChatGPT conversations. We finally got five primary categories of developers’ purposes (i.e., Task Delegation, Problem Resolution, Knowledge Acquisition, Solution Recommendation, and Concept Interpretation), and we analyzed their prevalence across the five sources. Figure 7 presents the distribution of the five purposes across five sources. Each circle is scaled to the total number of shared ChatGPT links for different sources, and the segments within each circle show the relative contribution of each purpose, making cross-category comparisons intuitive. For each category, we provide detailed descriptions of the results, followed by one example from our curated dataset. Detailed results are presented below.
Figure 7. Developers’ purposes of using ChatGPT links from the five sources
@@ 4.2. RQ2: Developers’ Purposes of Using ChatGPT Conversations
RQ2 Summary: Based on the shared developer-ChatGPT conversations, we identified and categorized developers’ purposes of using ChatGPT links during software development into five categories. The results show that Task Delegation is the dominant purpose (75-88% Code/Commits/Pull Requests). Problem Resolution (8-11% in Issues/Commits) and Concept Interpretation (¡6%) indicate GenAI’s context limitations. Knowledge Acquisition (21% in Discussions) and Solution Recommendation (17% in Discussions) thrive in collaborative ideation.
4.3. RQ3: Development-Related Activities in Shared ChatGPT Conversations
To address RQ3, which focuses on the development-related activities (coarse-grained level) discussed in shared ChatGPT conversations, we analyzed the distribution of shared ChatGPT links across the five data sources. Specifically, we applied the Constant Comparison method (Charmaz, 2014; Stol et al., 2016) to establish a categorization of the development-related activities, considering all extracted data items including filenames, path names, and other metadata (see Section 3.5). For example, requirements and user stories are categorized as “Requirement Analysis”; explicit change requests and regression fixes discussions are categorized as “Software Maintenance and Evolution”; and path names containing “xxx/testing” or mentions of unit tests are classified as “Software Testing”. Besides, software documentation includes various documents which can be integrated in lifecycle activities (e.g., requirements specifications, architecture and design documents, test reports). Considering that shared ChatGPT links also serve as a form of documentation recording developer-ChatGPT interactions during software development, we chose not to treat software documentation as a separate activity. Otherwise, all the shared ChatGPT links would fall under this documentation activity. The results of RQ3 reveal the breadth of developer interactions potentially facilitated by ChatGPT, as illustrated in Figures 8,  9, and 10.
@@ 4.3. RQ3: Development-Related Activities in Shared ChatGPT Conversations
Although there are various ways to categorize development-related activities, we considered and adapted the classification proposed in SWEBOK (Washizaki and others, 2024) and previous studies (Hou et al., 2024; Fan et al., 2023), which are widely recognized in the field of software engineering. Following this, we categorized the relevant activities involving shared developer-ChatGPT conversations, as shown in Figure 8. The development-related activities are defined in SWEBOK (Washizaki and others, 2024), and we grouped them into two primary categories: Lifecycle Activities and Supporting Activities. Regarding Lifecycle Activities, note that, activities from prior studies were not directly adopted. Instead, we identified lifecycle activities from prior studies (Washizaki and others, 2024; Hou et al., 2024; Fan et al., 2023), and retained only those consistently observed in our dataset (Li et al., 2025). Specifically, Requirements Analysis, Software Design, Software Deployment, Software Maintenance and Evolution, and Software Management are the categories from (Hou et al., 2024), and Software Implementation and Software Testing are the categories from (Fan et al., 2023).
In addition, based on the labeling results, we inductively abstracted two additional Supporting Activities (including Data Handling and Information Seeking), which do not map solely to a single SDLC phase but appear in ChatGPT-assisted multiple development phases. In this work, we used Development-Related Activities as an umbrella concept to encompass all activities evidenced through ChatGPT interactions in software development contexts.
  • Lifecycle Activities encompass seven general phases of the software development lifecycle (SDLC), that is, Requirements Analysis, Software Design, Software Implementation, Software Testing, Software Deployment, Software Maintenance and Evolution, and Software Management.
@@
RQ3 Summary: Analysis of shared ChatGPT links reveals that: (1) The number of shared ChatGPT links about the two lifecycle activities (i.e., Software Implementation and Software Maintenance & Evolution) far exceeds the number of other seven activities; (2) Software Implementation, Software Maintenance & Evolution are the dominant activities about which developers frequently share ChatGPT links.
4.4. RQ4: Software Development Tasks in Shared ChatGPT Conversations
To answer RQ4, we investigated the specific software engineering tasks (fine-grained level) within developer-ChatGPT interactions during development. Here, we adopted a broader view of tasks by conducting an in-depth analysis of developers’ intents, which encompass both action-oriented tasks (e.g., code generation) and knowledge-oriented goals (e.g., understanding specialized tools and techniques). While the latter may not constitute a task in the strict operational sense, we considered them equally relevant for characterizing developers’ practical use of ChatGPT, since seeking such knowledge is often an essential step toward completing a development task. Figure 11 presents a Sankey diagram that acts as a mapping framework linking GitHub data sources, development-related activities, and SE tasks. This decomposition visualization highlights the relationships between the five data sources, nine development-related activities, and 39 detailed SE tasks derived from ChatGPT interactions in our collected GitHub data, thereby reflecting how developers utilize ChatGPT in their daily workflows.
@@ 4.4. RQ4: Software Development Tasks in Shared ChatGPT Conversations
From a macroscopic perspective, our study reveals clear usage paths of ChatGPT on GitHub. For example, most frequently, developers delegate tasks to ChatGPT during the Software Implementation activity, accounting for the largest proportion of shared ChatGPT conversations. Within this activity, Code Generation and Completion emerges as the most prevalent task, indicating that developers primarily rely on ChatGPT as an implementation-level assistant for producing and completing source code.
@@
RQ4 Summary: We presented a mapping framework that links GitHub data sources, development-related activities, and SE tasks. Our analysis revealed 39 specific SE tasks across nine development-related activities, with Code Generation & Completion and Code Modification & Optimization emerging as the most prevalent SE tasks during software development. While code generation and maintenance tasks dominate, significant engagements were also observed in other crucial activities, highlighting ChatGPT’s broad utility as an assistant throughout the SE lifecycle.
@@ 5.1.1. Characteristics of ChatGPT Usage on GitHub
5.1.2. Developers’ Purposes of Using ChatGPT Conversations
@@ 5.2. Implications
For researchers, the findings point to promising opportunities to extend ChatGPT’s utility in certain areas, such as requirements engineering, by tailoring LLMs to domain-specific SE workflows. For example, the concentration of SE tasks in Software Implementation and Software Maintenance & Evolution implies that ChatGPT is primarily perceived as a coding assistant. Moreover, the findings highlight research opportunities to integrate ChatGPT in certain tasks (e.g., System Architecture Design), thereby improving its capability to boost productivity in underexploited areas. Besides, the study underscores ChatGPT’s transformative potential - not only in reshaping how developers engage in SE tasks, but also in redefining what constitutes a development-related activity in the AI era. This evolution calls for foundational knowledge in software engineering, such as SWEBOK (Washizaki and others, 2024), to adapt alongside emerging GenAI-driven software development practices.
@@
While this study provides valuable insights regarding the daily use of ChatGPT during software development, future research could delve deeper into the impact of ChatGPT on developer productivity. In particular, controlled experimental studies could be conducted to systematically compare development outcomes, such as task completion time, code quality, and defect rates, with and without ChatGPT assistance, thereby providing more rigorous evidence of its effectiveness.
For practitioners, understanding the impact of ChatGPT is crucial for strategic technology adoption during development. Although GenAI tools like ChatGPT exhibit great potential in enhancing developer efficiency (Vaillant et al., 2024), their limitations, such as the propensity to generate inaccurate or misleading information (often referred to as hallucination (Achiam et al., 2023)) or even unsafe code content (Chen et al., 2025), cannot be overlooked. Therefore, for researchers, this necessitates the collection of robust empirical evidence to assess the reliability and practical utility of ChatGPT in SE tasks. Future investigations should focus on determining the contexts in which these tools deliver substantial benefits and the contexts where they may introduce risks or inefficiencies (Yu et al., 2024). Addressing these concerns will be critical to ensuring the responsible and effective integration of LLMs into modern software development.
@@
Our results reveal that developers utilize ChatGPT in various development-related activities (particularly Software Implementation and Software Maintenance & Evolution). However, errors in ChatGPT’s outputs could potentially introduce inefficiencies or vulnerabilities in software systems (Liu et al., 2024a, b). Moreover, excessive cognitive reliance on GenAI tools might hinder the development of essential manual problem-solving skills, especially for junior developers. Given the potential of errors and biases in LLM-generated content (Chen et al., 2025; Li et al., 2024), it is imperative to establish robust safeguards, such as human oversight, rigorous validation processes, and strict adherence to ethical and legal standards. These measures are critical to maintaining the quality, security, and trustworthiness of AI-assisted development.